OpenStudy (anonymous):

given the function of f(x)= p-x / 2x-1 is undefined when x=h and p is constant . a) state the value of h b) given that 2 is mapped onto 1 through function f. find i) the values of p ii) the values if x if f(x) = 2x iii) the value of f^-1(-1)

OpenStudy (anonymous):

its undefined when denominator = 0 that is when h = 1/2

OpenStudy (anonymous):

can show me the way?

OpenStudy (anonymous):

substitute x = 1/2 into your function f(x)= p-x / 2x-1 when x = 1/2 f(x) =( p -1/2 ) / ( 2(1/2) - 1 ) =( p - 1/2) / 1 -1 = p - 1/2 / 0 anything divided by 0 is undefined